# BAP v2: An Enhanced Task Framework for Instruction Following in Minecraft Dialogues

### v2

@@ Acknowledgements.
We would like to thank Hetvi Patel, Sana Madhavan, Marc Canby, Risham Sidhu, Siddarth Madala, Rajarshi Haldar, Katya Yegorova, Adam Davies, Ansel Blume, and Sara Aghajanzadeh for their valuable help with the data annotation. This work was supported by Contract W911NF-15-1-0461 with the US Defense Advanced Research Projects Agency (DARPA) Communicating with Computers Program and the Army Research Office (ARO). Approved for Public Release, Distribution Unlimited. The views expressed are those of the authors and do not reflect the official policy or position of the Department of Defense or the U.S. Government. The work was also supported by the Distributed Research Experiences for Undergraduates (DREU) program, a joint project of the CRA Committee on the Status of Women in Computing Research (CRA-W) and the Coalition to Diversify Computing (CDC), which is funded in part by the NSF Broadening Participation in Computing program (NSF BPC-A #1246649). The work utilizes resources Kindratenko et al. (2020) supported by the National Science Foundation’s Major Research Instrumentation program, grant #1725729, as well as the University of Illinois at Urbana-Champaign.

### v3

@@ Acknowledgements.
    We would like to thank Hetvi Patel, Sana Madhavan, Marc Canby, Siddarth Madala, Rajarshi Haldar, Katya Yegorova, Adam Davies, Ansel Blume, and Sara Aghajanzadeh for their valuable help with the annotation, and Dan Roth, Martha Palmer, Janna Doppa and Sriraam Natarajan and all other members of our CwC team for many helpful discussions earlier in this project. This work was supported by Contract W911NF-15-1-0461 with the US Defense Advanced Research Projects Agency (DARPA) Communicating with Computers Program and the Army Research Office (ARO). Approved for Public Release, Distribution Unlimited. The views expressed are those of the authors and do not reflect the official policy or position of the Department of Defense or the U.S. Government. The work was also supported by the Distributed Research Experiences for Undergraduates (DREU) program, a joint project of the CRA Committee on the Status of Women in Computing Research (CRA-W) and the Coalition to Diversify Computing (CDC), which is funded in part by the NSF Broadening Participation in Computing program (NSF BPC-A #1246649). The work utilizes resources Kindratenko et al. (2020) supported by the National Science Foundation’s Major Research Instrumentation program, grant #1725729, as well as the University of Illinois at Urbana-Champaign, as well as the Delta advanced computing and data resource, which is supported by the National Science Foundation (award OAC 2005572) and the State of Illinois. Delta is a joint effort of the University of Illinois Urbana-Champaign and its National Center for Supercomputing Applications.